OpenStudy (anonymous):

4x - 3y = 17 5x + 4y = 60

OpenStudy (anonymous):

9x + y = 77

OpenStudy (anonymous):

\[4x-3y=17[1]\] \[5x + 4y = 60[2]\] \[[1]\times5\] \[5(4x-3y)=5(17)\] \[20x-15y=85[3]\] \[[2]\times4\] \[4(5x+4y)=4(60)\] \[20x+16y=240[4]\] \[[4]-[3]\] \[(20x+16y)-(20x-15y)=(240)-(85)\] \[16y+15y=155\] \[31y=155\] \[y=5\] To find the x-value, substitute the y-value (ie.5) into one of the four equations above (ie. [1], [2], [3] or [4])

OpenStudy (anonymous):

2x + 5y = 34 x + 2y = 14

OpenStudy (anonymous):

@dlee816 That's practically the same question. It just has different terms and values of x and y. Please use the method I provided you to solve it yourself. Giving questions that involve the same method as before wastes other people's time when there are many others who need their questions to be answered.

OpenStudy (anonymous):

or assisted*